# Mitigating the Linguistic Gap with Phonemic Representations for Robust Multilingual Language Understanding

### v1

@@ 4 Results
Let $\mathcal{D}$ denote a domain as a distribution over text feature input $\mathcal{X}$, such as the sequence of word embeddings or one-hot vectors, and a labeling function $f:\mathcal{X}\rightarrow\{0,1\}$. Assuming a binary classification task, our goal is to learn a hypothesis $h:\mathcal{X}\rightarrow\{0,1\}$ that is expected to minimize a risk $\varepsilon_{D}(h,f):=\mathbb{E}_{x\sim\mathcal{D}}[\mathbb{I}(f(x)\neq h(x))]$ and has a small risk-deviation over two domains $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$. Then, to formalize the cross-lingual performance gap, we first need a discrepancy measure between two languages. By following Ben-David et al. (2010), we adopt $\mathcal{H}$-divergence (See Appendix B for its definition) to quantify the distance between two language distributions.
@@
Let $h:\mathcal{X}\rightarrow[0,1]$ be a real-valued function in a hypothesis class $\mathcal{H}$ with a pseudo dimension $\mathcal{P}dim(\mathcal{H})=d$. If $\hat{\mathcal{D}}_{A}$ and $\hat{\mathcal{D}}_{B}$ are the empirical distribution constructed by $n$-size i.i.d. samples, drawn from $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ respectively, then for any $\delta\in(0,1)$, and for all $h$, the bound below hold with probability at least $1-\delta$.
$|\varepsilon_{\mathcal{D}_{A}}(h,f)-\varepsilon_{\mathcal{D}_{B}}(h,f)|\leq{1\over 2}d_{\mathcal{H}\Delta\mathcal{H}}(\hat{\mathcal{D}}_{A},\hat{\mathcal{D}}_{B})$
$+2\sqrt{{{d\log(2n)+\log(2/\delta)}\over n}}$
where $\mathcal{H}\Delta\mathcal{H}:=\{h(x)\oplus h^{\prime}(x)|h,h^{\prime}\in\mathcal{H}\}$ given $\oplus$ as a xor operation (proof is in Appendix B). We see that performance gap between two lanauges is bounded from above with a distribution divergence plus an irreducible term defined by problem setup. That is, if we reduce the divergence between language distributions, the expected performance gap can also be reduced accordingly.
To investigate whether this is indeed a case or not, we provided embedding space similarity and logit-space Sinkhorn distance between different languages in Figure 3. We argue that phonemic representation’s relatively mild performance gap is achieved by reducing linguistic gaps which is confirmed in the embedding space (high CKA) and final output space (low Sinkhorn distance).
@@ 5 Conclusion
Towards robust multilingual language modeling, we argue that mitigating the linguistic gap between different languages is crucial. Moreover, we advocate the use of IPA phonetic symbols as a universal language representation that partially bridges such linguistic gaps without any complicated cross-lingual training phase. Empirical validation on three representative NLP tasks demonstrates the superiority of phonemic representation compared to subword and character-based language representation in terms of the cross-lingual performance gap and linguistic gap. Theoretical analysis of the cross-lingual performance gap explains such promising results of phonemic representation.
@@ 6 Limitations
While we have shown that phonemic representation induces a small cross-lingual linguistic gap, therefore a small performance gap, the absolute performance of this phonemic representation is still lacking compared to subword-level models. We spur the necessity of putting research attention to developing phoneme-based LMs. Moreover, there is no such large phonemic language model beyond the BERT-base-size architecture, so we confine the scope of our empirical validation to BERT-base-size LMs. Thus, we can not ensure the effectiveness of IPA representation when adopted within modern large language models, such as LLaMa 2 Touvron et al. (2023). Additionally, we performed evaluation with a limited languages (up to 8), so it is unclear whether IPA language representations are effective for other numerous languages (especially low-resource ones) or not.
@@ 7 Ethics Statement
Appendix A Details on Experiment Setup
A.1 Models
We compare LMs with three different types of language representation: subword, character, and phoneme. All models555pre-trained weights were obtained from https://huggingface.co/models are pre-trained on multilingual data that covers around 100 languages from Wikipedia dump files. We employ an off-the-shelf MLLMs, mBERT Devlin et al. (2018). for a subword-based model which is trained on 104 languages with byte-pair encoding (vocabulary size of 110k). The model is pre-trained with the objective of masked language modeling and next sentence prediction tasks. For a character-based model, we utilize CANINE Clark et al. (2022), which is a tokenization-free LM that directly maps each character to its codepoint by hashing. It is pre-trained on the same data and training objectives as mBERT. This prevents unknown tokens, enabling the model to handle a large amount of distinct characters. CANINE is pre-trained on the same data and training objectives as mBERT. For a phoneme-based model, we adopt XPhoneBERT Nguyen et al. (2023), which has the same model architecture as mBERT and it is trained on 94 languages with learning objective of dynamic masked language modeling.
While character-level models are known to better generalize to low-resource languages Clark et al. (2022), their general performance falls behind subword-based models. For a fair comparison between the representations, we compare phoneme-based model to character-based model instead of directly comparing it to widely used subword-based model (i.e., mBERT), and leave further improvements of overall performance as future work.
A.2 Data
G2P Conversion.
In order to pass the input to phoneme-based model, we employed G2P (Grapheme-to-Phoneme) conversion on the data to obtain an IPA version of the input. This conversion was done with Epitran [^6] https://github.com/dmort27/epitran Mortensen et al. (2018) , an external tool for G2P conversion. After converting to IPA, we insert white space between every character to make it compatible with the XPhoneBERT tokenizer.
Languages.
We categorize languages into three groups—high-resource (eng, deu, fra), low-resource (urd, hin, swa), and typologically distinct (kor, ukr)—to analyze the impact of phonemic representation with respect to each category. For high/low-resource language, we refer to Wu and Dredze (2020) and treat the languages with a wikisize under 8 as low-resource languages, and those with wikisize above 11 high-resource languages. On the other hand, typologically distinct languages are chosen with reference to English (they are orthographically and typologically different) [^7] Note that this categorization is not mutually exclusive and that Urdu and Hindi can also be considered as typologically distant from English..
A.3 Downstream Tasks
We adopt the cross-lingual generalization benchmark tasks suggested in XTREME Hu et al. (2020). Among them, we selected two types of tasks; classification and structured prediction.
Sentence-level Classification.
XTREME supports some sentence-level classification tasks, requiring semantic understanding of given sentences to make a prediction. We employ the XNLI Conneau et al. (2018) corpus, which is well-known for cross-lingual evaluation, to train and evaluate the model on different languages.
Token-level Classification.
Structured prediction tasks from Hu et al. (2020) include POS tagging and NER. Both tasks require labeling each token from the model. These types of tasks were previously analyzed as being relatively independent from the data size of each language used for pre-training Hu et al. (2020). We find this particularly suitable in our scenario where two models with different pre-training strategy are compared. We utilize the corpora from Universal Dependencies [^8] https://universaldependencies.org/ , v2.13, 148 languages, released Nov 15, 2023. for POS tagging, and WikiAnn Pan et al. (2017) with train, dev, test splits following Rahimi et al. (2019) for NER. While all 8 languages are employed for NER, we do not consider swa for POS tagging since Universal Dependencies does not provide a Swahili treebank.
Evaluation.
For evaluation, we follow the common practice of each task, evaluating XNLI task with top 1 accuracy and NER, POS tagging with F1 score. All metrics used are computed via functions from scikit-learn  [^9] https://scikit-learn.org/stable/index.html python library.
A.4 Implementation details.
We fine-tuned all models with AdamW Loshchilov and Hutter (2018) optimization setting batch size as 128 over 20 epochs for XNLI and 40 epochs for NER and POS Tagging tasks. Here, we adopt a cosine learning rate scheduler with a warm-up on the XNLI task. For all models, we adopt mixed precision training Micikevicius et al. (2018) provided by PyTorch, i.e., with autocast() loop for saving computational cost and energy consumption. For XNLI task, we used the embedding of [CLS] token for mBERT, and used the mean of all tokens except [CLS] token for CANINE and XPhoneBERT. All experiments were done with the random seed fixed to 42.
A.5 Number of parameters of the models.
Model Num. of params
mBERT-base 177,853,440
CANINE-c 132,082,944
XPhoneBERT 87,554,304
Table 3: The number of parameters of each model.
A.6 Computational resources.
All experiments on POS tagging were done on a single NVIDIA GeForce RTX 3090, and those on XNLI and NER were done on a single NVIDIA A6000. Total GPU hours for all experiments are 29 days.
A.7 Hyperparameter sweep.
@@ A.3 Downstream Tasks
NER learning rate [1e-5, 5e-5, 1e-4] 1e-4 5e-5 5e-5
weight decay [1e-4, 1e-3, 1e-2] 1e-3 1e-4 1e-2
POS learning rate [1e-5, 5e-5, 1e-4, 3e-4, 1e-2] 1e-4 3e-4 1e-4
weight decay [1e-4, 1e-3, 1e-2] 1e-4 1e-2 1e-2
@@
A.8 Datasets, Statistics, and License.
In Table 5, we provide the datasets, their statistics, and license.
Dataset Lang. Train Dev Test License
XNLI eng 393k 2.49k 5.01k CC BY-NC-4.0
swa
urd
WikiAnn eng 20k 10k 10k ODC-BY
deu 20k 10k 10k
fra 20k 10k 10k
kor 20k 10k 10k
ukr 20k 10k 10k
urd 20k 1k 1k
swa 1k 1k 1k
hin 5k 1k 1k
UD eng 12.5k 2k 2k CC BY-SA 4.0
deu 13.8k 0.8k 1k CC BY-SA 4.0
fra 14.5k 1.5k 0.4k CC BY-SA 4.0
kor 23k 2k 2.3k CC BY-SA 4.0
ukr 5.5k 0.7k 0.9k CC BY-NC-SA 4.0
urd 4k 0.6k 0.5k CC BY-NC-SA 4.0
Table 5: Statistics and license types for datasets. The table lists the number of examples in the training, development, and testing sets for languages in the XNLI, WikiAnn, and UD datasets. It specifies the licensing conditions: CC-BY permits sharing and adaptation, CC BY-NC is for non-commercial usage, CC BY-SA mandates share-alike distributions, and CC BY-NC-SA combines non-commercial with share-alike terms. All datasets are strictly used within the bounds of these licenses.
Appendix B Details on Theoreoretical Analysis
We aim to diminish the performance gap between different languages by adopting IPA as a universal language representation. Motivated by domain adaptation literature Kifer et al. (2004); Ben-David et al. (2010), we present a theoretical justification of IPA for robust multilingual modeling by providing a bound for cross-lingual performance gap.
Let $\mathcal{D}$ denote a domain as a distribution over text feature input $\mathcal{X}$, such as the sequence of word embeddings or one-hot vectors, and a labeling function $f:\mathcal{X}\rightarrow\{0,1\}$. Assuming a binary classification task, our goal is to learn a hypothesis $h:\mathcal{X}\rightarrow\{0,1\}$ that is expected to minimize a risk $\varepsilon_{D}(h,f):=\mathbb{E}_{x\sim\mathcal{D}}[\mathbb{I}(f(x)\neq h(x))]$ and has a small risk-deviation over two domains $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$. Then, to formalize the cross-lingual performance gap, we first need a discrepancy measure between two languages. By following Ben-David et al. (2010), we adopt $\mathcal{H}$-divergence to quantify the distance between two language distributions.
Definition B.1 ($\mathcal{H}$-divergence; Ben-David et al. (2006)).
@@ Appendix B Details on Theoreoretical Analysis
Theorem B.2.
Let $h:\mathcal{X}\rightarrow[0,1]$ be a real-valued function in a hypothesis class $\mathcal{H}$ with a pseudo dimension $\mathcal{P}dim(\mathcal{H})=d$. If $\hat{\mathcal{D}}_{A}$ and $\hat{\mathcal{D}}_{B}$ are the empirical distribution constructed by $n$-size i.i.d. samples, drawn from $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ respectively, then for any $\delta\in(0,1)$, and for all $h$, the bound below hold with probability at least $1-\delta$.
$|\varepsilon_{\mathcal{D}_{A}}(h,f)-\varepsilon_{\mathcal{D}_{B}}(h,f)|\leq{1\over 2}d_{\mathcal{H}\Delta\mathcal{H}}(\hat{\mathcal{D}}_{A},\hat{\mathcal{D}}_{B})$
$+2\sqrt{{{d\log(2n)+\log(2/\delta)}\over n}}$
@@
Finally, by plugging the Lemma B.3 into the above bound, we have Theorem B.2.
From Theorem B.2, we see that the difference between true risks across language domains is bounded by an empirical estimation of the divergence ($d_{\mathcal{H}\Delta\mathcal{H}}$) between those two domains plus an irreducible term defined by problem setup. Thus, if we reduce the divergence between language distributions, the expected performance gap can also be reduced accordingly. To investigate whether this is indeed a case or not, we provided the embedding-space similarity and logit-space Sinkhorn distance between different languages in Figure 3. We argue that phonemic representation’s relatively mild performance gap is achieved by reducing linguistic gaps in the embedding space (high CKA) and final output space (low Sinkhorn distance) those are the proxy of $\mathcal{H}$-divergence.

### v3

@@ 4.4 Connecting Performance Gap and Linguistic Gap
Let $\mathcal{D}$ denote a domain as a distribution over text feature input $\mathcal{X}$, such as the sequence of word embeddings or one-hot vectors, and a labeling function $f:\mathcal{X}\rightarrow[0,1]$. Assuming a binary classification task, our goal is to learn a hypothesis $h:\mathcal{X}\rightarrow\{0,1\}$ that is expected to minimize a risk $\varepsilon_{D}(h,f):=\mathbb{E}_{x\sim\mathcal{D}}[|f(x)-h(x)|]$ and has a small risk-deviation over two domains $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$. Then, to formalize the cross-lingual performance gap, we first need a discrepancy measure between two languages. By following Ben-David et al. (2010), we adopt $\mathcal{H}$-divergence (See Appendix C for its definition) to quantify the distance between two language distributions.
@@
Let $h:\mathcal{X}\rightarrow\{0,1\}$ be a function in a hypothesis class $\mathcal{H}$ with a pseudo dimension $\mathcal{P}dim(\mathcal{H})=d$. If $\hat{\mathcal{D}}_{A}$ and $\hat{\mathcal{D}}_{B}$ are the empirical distribution constructed by $n$-size i.i.d. samples, drawn from $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ respectively, then for any $\delta\in(0,1)$, and for all $h$, the bound below hold with probability at least $1-\delta$.
$|\varepsilon_{\mathcal{D}_{A}}(h,f)-\varepsilon_{\mathcal{D}_{B}}(h,f)|\leq\frac{1}{2}d_{\mathcal{H}\Delta\mathcal{H}}(\hat{\mathcal{D}}_{A},\hat{\mathcal{D}}_{B})$
$+2\sqrt{\frac{d\log(2n)+\log(2/\delta)}{n}}$
where $\mathcal{H}\Delta\mathcal{H}:=\{h(x)\oplus h^{\prime}(x)|h,h^{\prime}\in\mathcal{H}\}$ given $\oplus$ as a xor operation (proof is in Appendix C). We see that performance gap between two languages is bounded from above with a distribution divergence plus an irreducible term defined by problem setup. That is, if we reduce the divergence between language distributions, the expected performance gap can also be reduced accordingly.
To investigate whether this is indeed a case or not, we provided embedding space similarity and logit-space Sinkhorn distance Cuturi (2013) between different languages in Figure 3. We argue that phonemic representation’s relatively mild performance gap is achieved by reducing linguistic gaps which is confirmed in the embedding space (high CKA) and final output space (low Sinkhorn distance).
@@ 5 Conclusion
Towards robust multilingual language modeling, we argue that mitigating the linguistic gap between different languages is crucial. Moreover, we advocate the use of IPA phonetic symbols as a universal language representation partially bridges such linguistic gaps without any complicated cross-lingual training phase. Empirical validation on three representative NLP tasks demonstrates the superiority of phonemic representation compared to subword and character-based language representation in terms of the cross-lingual performance gap and linguistic gap. Theoretical analysis of the cross-lingual performance gap explains such promising results of phonemic representation.
@@ 6 Limitations
While we have shown that phonemic representation induces a small cross-lingual linguistic gap, therefore a small performance gap, the absolute performance of this phonemic representation is still lacking compared to subword-level models. We spur the necessity of putting research attention to developing phoneme-based LMs. Moreover, there is no such large phonemic language model beyond the BERT-base-size architecture, so we confine the scope of our empirical validation to BERT-base-size LMs. This also means the experiments rely on existing pre-trained models, limiting control over their pre-training settings. Since the models were trained on different language sets and pre-training objectives (as noted in 3.1), it is important to verify these findings in a controlled environment. Additionally, we performed evaluation with a limited languages (up to 12), so it is unclear whether IPA language representations are effective for other numerous languages (especially low-resource ones) or not.
7 Impact Statement
Ensuring strong multilingual understanding capability in our language modeling systems is increasingly crucial given the openness of frontier-level AI produce Dubey et al. (2024). Enhanced cross-lingual transferability by leveraging our findings may have some positive impacts on diverse domains such as healthcare He et al. (2023); Ryu et al. (2024) and education Han et al. (2023); Oh et al. (2024) which are very close to our daily life as well as on further development of cutting-edge multimodal large language models Kim et al. (2023); Yin et al. (2023).
@@ 8 Ethics Statement
Acknowledgements
We sincerely thank Haneul Yoo at KAIST who provided constructive feedback on the manuscript. This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2022-00143911, AI Excellence Global Innovative Leader Education Program)
Appendix A Dataset Statistics
In Table 4, we provide the dataset statistics. For the experiments, we used train set for training and validation set for evaluation.
Dataset Lang. Train Dev Test
FLORES+ eng - 1.2k -
fra
rus
ita
hun
ukr
kor
tur
fin
hin
XNLI eng 393k 2.49k 5.01k
swa
urd
WikiAnn eng 20k 10k 10k
fra 20k 10k 10k
rus 20k 10k 10k
ita 20k 10k 10k
hun 20k 10k 10k
ukr 20k 10k 10k
kor 20k 10k 10k
tur 20k 10k 10k
fin 20k 10k 10k
hin 5k 1k 1k
UD eng 12.5k 2k 2k
fra 14.5k 1.5k 0.4k
rus 16k 0.9k 0.9k
ita 13k 0.6k 0.5k
hun 0.9k 0.4k 0.4k
ukr 5.5k 0.7k 0.9k
kor 23k 2k 2.3k
tur 15k 1.6k 1.6k
fin 12k 1.4k 1.6k
hin 13k 1.7k 1.7k
Table 4: Dataset statistics for datasets used in experiments: FLORES+, XNLI, WikiAnn, Universal Dependencies Tree Bank. For FLORES+ dataset, we used devtest set with 1,012 sentences.
Appendix B Hyperparameter sweep.
@@ Appendix A Dataset Statistics
NER learning rate [3e-5, 5e-5, 1e-4, 3e-4] - 5e-5 (en, fr, it, hu, ko, tr), 1e-4 (ru, uk, fi, hi) 3e-5 (ru, it), 5e-5 (en, fr, hu, uk, tr, fi, hi), 1e-4 (ko)
weight decay 1e-2 - 1e-2 1e-2
POS learning rate [3e-5, 5e-5, 1e-4, 3e-4] - 5e-5 (ru, uk, tr), 1e-4 (en, fr, fi, hi), 3e-4 (it, hu, ko) 5e-5 (en), 1e-4 (fr, ru, it, hu, uk, ko, tr, fi, hi)
weight decay 1e-2 - 1e-2 1e-2
@@
Appendix C Details on Theoreoretical Analysis
We aim to diminish the performance gap between different languages by adopting IPA as a universal language representation. Motivated by domain adaptation literature Kifer et al. (2004); Ben-David et al. (2010); Oh et al. (2023), we present a theoretical justification of IPA for robust multilingual modeling by providing a bound for cross-lingual performance gap.
Let $\mathcal{D}$ denote a domain as a distribution over text feature input $\mathcal{X}$, such as the sequence of word embeddings or one-hot vectors, and a labeling function $f:\mathcal{X}\rightarrow[0,1]$. Assuming a binary classification task, our goal is to learn a hypothesis $h:\mathcal{X}\rightarrow\{0,1\}$ that is expected to minimize a risk $\varepsilon_{D}(h,f):=\mathbb{E}_{x\sim\mathcal{D}}[|f(x)-h(x)|]$ and has a small risk-deviation over two domains $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$. Then, to formalize the cross-lingual performance gap, we first need a discrepancy measure between two languages. By following Ben-David et al. (2010), we adopt $\mathcal{H}$-divergence to quantify the distance between two language distributions.
Definition C.1 ($\mathcal{H}$-divergence; Ben-David et al. (2006)).
@@ Appendix C Details on Theoreoretical Analysis
Theorem C.2.
Let $h:\mathcal{X}\rightarrow\{0,1\}$ be a function in a hypothesis class $\mathcal{H}$ with a pseudo dimension $\mathcal{P}dim(\mathcal{H})=d$. If $\hat{\mathcal{D}}_{A}$ and $\hat{\mathcal{D}}_{B}$ are the empirical distribution constructed by $n$-size i.i.d. samples, drawn from $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ respectively, then for any $\delta\in(0,1)$, and for all $h$, the bound below hold with probability at least $1-\delta$.
$|\varepsilon_{\mathcal{D}_{A}}(h,f)-\varepsilon_{\mathcal{D}_{B}}(h,f)|\leq\frac{1}{2}d_{\mathcal{H}\Delta\mathcal{H}}(\hat{\mathcal{D}}_{A},\hat{\mathcal{D}}_{B})$
$+2\sqrt{\frac{d\log(2n)+\log(2/\delta)}{n}}$
@@
Finally, by plugging the Lemma C.3 into the above bound, we have Theorem C.2.
From Theorem C.2, we see that the difference between true risks across language domains is bounded by an empirical estimation of the divergence ($d_{\mathcal{H}\Delta\mathcal{H}}$) between those two domains plus an irreducible term defined by problem setup. Thus, if we reduce the divergence between language distributions, the expected performance gap can also be reduced accordingly. To investigate whether this is indeed a case or not, we provided the embedding-space similarity and logit-space Sinkhorn distance between different languages in Figure 3. We argue that phonemic representation’s relatively mild performance gap is achieved by reducing linguistic gaps in the embedding space (high CKA) and final output space (low Sinkhorn distance) those are the proxy of $\mathcal{H}$-divergence.